The sum of three different integers can never be zero.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

The given statement is False.
For Example:

Consider 3 different integers 5, 10 and -15.
Sum of 3 integers = 5 + 10 + (-15)

= 5 + 10–15 = 15–15
=0

Therefore, the sum of three different integers can be zero.

Thanks